OpenStudy (anonymous):

factor: 6x^2 - x - 2

OpenStudy (anonymous):

hm...it might be best if you use the quadratic formula for this as the solutions are not obvious. The solutions are real numbers but its hard to get it from shortcut method.

OpenStudy (anonymous):

when you get the solutions, just equate them to x. like one of the solutions is -1/2: x = -1/2 x + (1/2) = 0 (x + (1/2))(x + (the other solution)) = 0